Combustion of hydrocarbons such as decane ( C 10 H 22 ) produces carbon dioxide, a "greenhouse gas." Greenhouse gases in the Earth's atmosphere can trap the Sun's heat, raising the average temperature of the Earth. For this reason there has been a great deal of international discussion about whether to regulate the production of carbon dioxide. 1. Write a balanced chemical equation, including physical state symbols, for the combustion of liquid decane into gaseous carbon dioxide and gaseous water. 2. Suppose 0.180 kg of decane are burned in air at a pressure of exactly 1 atm and a temperature of 11.0 °C . Calculate the volume of carbon dioxide gas that is produced. Be sure your answer has the correct number of significant digits.
Answer #1

1.

2C10H22(l) + 31O2(g) --------------> 20CO2(g) + 22H2O(g)

2.

2C10H22(l) + 31O2(g) --------------> 20CO2(g) + 22H2O(g)

2 moles of C10H22 combustion to gives 20 moles of CO2

2*142g of C10H22 combustion to gives 20 moles of CO2

180g of C10H22 combustion to gives = 20mole*180g/(2*142g)   = 12.676 moles of CO2

n = 12.676moles

T   = 11+273 = 284K

P = 1atm

PV = nRT

V = nRT/P

     = 12.676*0.0821*284/1   = 295.56L of CO2

the volume of CO2 = 295.56L of CO2